Which number is equivalent to 2/5?
A) 2 and 1/2
B) 0.4
C) 5/2
D) 2 and 1/5
Thanks

Answer by jmg(22)

If you have multiple choice problems I always work the problems by eliminating wrong answers.
You should automatically be able to eliminate b. You should know that 2/5 is not equal to 5/2.
You should also be able to automatically eliminate A and D. Both of those numbers have a whole number of 2 attached to them.
You should know that 2/5 will never be bigger than 1. If you only have 2/5 then you definitely don't have 5/5, which would equal 1, so you definitely don't have 2.
That leaves answer choice B.
You can show that 2/5 = 0.4 in one of two ways:
Calculator (if you are allowed to use one): Divide the top
number by the bottom number. 2/5 = 0.4
If you can't use a calculator. Then change 2/5 to a number out of 100.
2/5 = 40/100
40/100 = 40% = 0.4 or 0.40
Hope this helps.